# ERX: A Fast Real-Time Anomaly Detection Algorithm for Hyperspectral Line-Scanning

### v1

@@ V-B Limitations & Future Work
Creating an open-source Python repository improves accessibility and reproducibility for future projects, but at the cost of slower processing speeds. Implementing these algorithms in faster languages like C++ would significantly improve their processing speeds. If faster languages were used, more algorithm types, such as those using kernel and window methods, may become feasible for real-time applications. Performing field tests of these algorithms for extended periods, beyond simulations, is crucial for validating the results, adaptability, and practicality of the ERX algorithm.
@@
This work lays a foundation for real-time RX-based pixel anomaly detection, although the generally low F1 and AUC scores highlighted an ongoing challenge in detecting anomalies at the pixel level. Grouping pixels into anomalous objects and locating them is a more useful task in practice and may yield better results with object-based metrics. Future research could explore unsupervised, self-supervised, and synthetically trained object detection algorithms (e.g., YOLO [41]). Rather than line-wise detection, using batches of hyperspectral lines could be more practical for detecting and locating anomalous objects in real-time, or at least near real-time.
@@ VI Conclusion
This research introduced the Exponentially moving RX (ERX) algorithm for real-time anomaly detection in hyperspectral line-scanning. ERX is fast and scalable to the high dimensions of hyperspectral imagery. It adapts quickly to changing scenery and is robust in detecting anomalies using less corrected data (radiance instead of reflectance).
Tuning ERX’s momentum parameter (exponential decay) improves detection performance based on the anomaly frequency and the variation of the scenery. Three datasets were introduced, and four other algorithms were implemented onboard a Jetson Xavier NX compute module to evaluate ERX’s performance. ERX excelled in detection and speed on the beach dataset, being the only algorithm fast enough to run in real-time. ERX had the best detection performance on the synthetic dataset, and it was the fastest algorithm on the SNP dataset.
Overall, ERX had the strongest combined performance, but the datasets were challenging for all algorithms. This highlights the need for new anomaly detection methods. There are promising pathways for future work in anomaly detection for hyperspectral line-scanning in real-time (or near real-time); particularly with unsupervised and self-supervised algorithms that group anomalous pixels into objects and locate them, adaptive or automatic threshold selection, and field tests that validate the algorithms’ practicality.
Acknowledgment

### v2

@@ V-D Limitations & Future Work
Creating an open-source Python repository improves accessibility and reproducibility for future projects, but at the cost of slower processing speeds. Implementing these algorithms in faster languages like C++ would significantly improve their processing speeds. If faster languages were used, more algorithm types, such as those using kernel and window methods, may become feasible for real-time applications. Performing field tests for extended periods, beyond simulations, is crucial to validating the practicality of the ERX algorithm.
@@
Despite ERX’s overall success, the challenges in separating anomalies from the background in complex datasets remain a persistent issue in pixel-level detection. Grouping pixels into anomalous objects and locating them is a more useful task in practice and may yield better results with object-based metrics. Future research could investigate unsupervised, self-supervised, or synthetically trained object detection algorithms (e.g., YOLO [83]) to detect anomalies in any scene containing geometric distortions. Rather than line-by-line detection, using batches of hyperspectral lines could be more practical for detecting and locating anomalous objects in real time, or at least near real time.
@@ VI Conclusion
This research introduced the Exponentially moving RX (ERX) algorithm for real-time anomaly detection in hyperspectral line scanning. ERX is fast and scalable to the high dimensions of hyperspectral imagery. It adapts quickly to changing scenery and is robust in detecting anomalies using HSIs that have not been geometrically corrected. Three larger and more complex datasets were introduced to better reflect the challenges of hyperspectral anomaly detection for line scanning, and four other algorithms were implemented onboard a Jetson Xavier NX compute module to evaluate ERX’s performance with limited compute power.
ERX excelled in detection and speed on the beach and synthetic datasets, being the only algorithm fast enough to run in real-time. ERX was the second fastest algorithm on the SNP dataset, and still maintained competitive performance. Sparse random projection effectively reduced dimensionality, enhancing both speed and detection. Maximum benefits are obtained at low dimensions, but excessively low dimensions increase model variance due to inherent randomness. Adjusting ERX’s momentum parameter enhances detection performance according to anomaly frequency and scenery variation, while the exponentially moving mean and covariance ensure adaptability and faster processing speeds.
ERX performed best overall, yet the datasets were challenging and open up some promising pathways for future work. Some opportunities include field tests to validate algorithm practicality, adaptive and automatic threshold selection, and unsupervised and self-supervised algorithms that group anomalous pixels into objects and locate them.
Acknowledgments